# Reinforcement Learning via Self-Distillation

### v1

@@ 1 Introduction
Nevertheless, the dominant RL recipe for LLM post-training remains bottlenecked by credit assignment. Most current approaches operate in the setting of reinforcement learning with verifiable rewards (RLVR): given a question $x$, the model samples an answer $y\sim\pi_{\theta}(\cdot\mid x)$} and receives a scalar reward $r\in\mathbb{R}$, often binary (e.g., unit-tests pass/fail in code generation). Modern policy gradient RLVR methods such as Group Relative Policy Optimization (GRPO; Shao et al., 2024) estimate advantages from these sparse outcome rewards. Furthermore, when all rollouts in a group receive the same (often zero) reward, GRPO advantages collapse to zero and learning stalls. To overcome this sparsity, one might prefer distillation from a strong teacher (Guo et al., 2025; Yang et al., 2025; Lu & Thinking Machines Lab, 2025; Guha et al., 2026), which provides dense, token-level supervision. However, strong teachers are often unavailable in online learning, where the goal is to raise the capability ceiling beyond existing models.
@@
  • Learning without rich feedback (§3): We evaluate standard RLVR environments that do not return any feedback beyond scalar rewards. Here, SDPO treats successful attempts sampled in the current batch as “feedback” for failed attempts on the same question. We perform training runs on scientific reasoning and tool use, starting with Qwen3-8B and Olmo3-7B-Instruct. We find that SDPO outperforms a strong GRPO baseline that integrates recent improvements: 68.8% vs. 64.1% final accuracy on aggregate. SDPO achieves higher accuracy with up to $7\times$ shorter generation lengths compared to GRPO, demonstrating that effective reasoning need not be verbose.
@@ 2 SDPO: Self-Distillation Policy Optimization
We can derive the SDPO gradient as follows (see Section˜B.1 for details):
@@
Let $\mathcal{V}$ be the set of tokens in the vocabulary. The gradient of $\mathcal{L}_{\mathrm{SDPO}}$ is
$\boldsymbol{\nabla}_{\!\!\theta}\,\mathcal{L}_{\mathrm{SDPO}}(\theta)=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\sum_{t=1}^{|y|}\sum_{\hat{y}_{t}\in\mathcal{V}}\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\cdot\log\frac{\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})}{\pi_{\theta}(\hat{y}_{t}\mid x,f,y_{<t})}\right].\vskip 4.30554pt$ (2)
@@ 2.1 Comparison to RLVR
Note that the SDPO gradient is a (negated) logit-level policy gradient where the advantages are estimated using the self-teacher. [^2] See Section˜A.3 for a detailed comparison of the SDPO gradient to the standard policy gradient. We can therefore reuse standard RLVR implementations and simply swap out the advantages. Let $y_{i}$ be the $i$-th rollout from a rollout group of size $G$ for question $x$, then we have:
$A_{i,t}^{\mathrm{GRPO}}(\hat{y}_{i,t}):=\mathbbm{1}\{{y_{i,t}=\hat{y}_{i,t}}\}\left(r_{i}-\mathrm{mean}\{r_{i}\}_{i=1}^{G}\right),\quad A_{i,t}^{\mathrm{SDPO}}(\hat{y}_{i,t})=\log\frac{\pi_{\theta}(\hat{y}_{i,t}\mid x,f_{i},y_{i,<t})}{\pi_{\theta}(\hat{y}_{i,t}\mid x,y_{i,<t})}.$
The GRPO advantages are zero on any non-generated token and constant within a rollout $y_{i}$. [^3] We use the GRPO (Shao et al., 2024) advantage without normalization (Liu et al., 2025b). In contrast, the SDPO advantages are zero only for tokens where student and teacher perfectly agree. The SDPO advantage is positive for tokens which are more likely under the teacher while being negative for tokens which are less likely under the teacher. Thus, SDPO can be seen as a direct extension of standard RLVR methods in two ways:
@@
This tight connection to RLVR methods also enables a straightforward extension of the SDPO gradient from Equation˜2 to off-policy data via PPO-style clipped importance sampling (Schulman et al., 2017), see Section˜A.3.
@@ 2.2 Compute time & memory
Naively computing the KL divergence between student and teacher requires holding full logits of both models in memory. To avoid this, we approximate the KL divergence in the SDPO loss by performing top-$K$ distillation (i.e., only computing the top-$K$ logits of the student and the corresponding logits of the teacher alongside a term capturing the tail probability; cf. Section˜A.2). With a reasonable choice of $K$ (e.g., K=100{K=100}), this avoids virtually any memory overhead while capturing most of the information.
@@ 2.3 Stability improvements
We find that two practical modifications significantly enhance the training stability of SDPO. First, we employ a regularized self-teacher, implemented either via an exponential moving average (EMA) of the student parameters or by interpolating the current teacher with the initial teacher (cf. Section˜A.1). As detailed later, both strategies effectively stabilize learning. Second, we adopt the symmetric Jensen-Shannon divergence for the distillation loss; this formulation has similarly been shown to improve stability in on-policy distillation from external teachers (Agarwal et al., 2024).
@@ 3 Learning without Rich Environment Feedback
Figure 6: Training progression of Olmo3-7B-Instruct on Chemistry. We report the average accuracy across 16 samples per question and a rolling average of response lengths over 5 steps. We report GRPO with the optimal hyperparameters for this model and task.
@@ 3.1 Experimental setting
  • Science Q&A (Chemistry, Physics, Biology, Materials science): Undergraduate-level scientific reasoning using reasoning subsets (L3) from SciKnowEval (Feng et al., 2024).
@@
We perform a train-test split to test in-domain generalization. We use Qwen3-8B (Yang et al., 2025) and Olmo3-7B-Instruct (Olmo et al., 2025) as initial checkpoints and report avg@16 relative to wall-clock training time, excluding initialization & validation.
@@
We compare SDPO to an improved variant of GRPO (Shao et al., 2024), which incorporates several recent modifications (Olmo et al., 2025; Khatri et al., 2026) such as asymmetric clipping (Yu et al., 2025), avoiding biased normalization (Liu et al., 2025b), and correcting for off-policy data when using efficient inference frameworks (Yao et al., 2025). We integrate these modifications into a GRPO implementation that represents a strong baseline, as detailed in Equation˜10 in Section˜A.3. GRPO enables off-policy training through PPO’s clipped importance weighting (Schulman et al., 2017). We additionally report the special case of on-policy GRPO (matching the hyperparameters of vanilla SDPO). For both baselines, we perform a hyperparameter sweep and report results for the models that achieve the highest validation performance across all target tasks. Hyperparameters and training details are provided in Appendix˜E. We use the verl library (Sheng et al., 2025) for fast multi-GPU training.
@@ 3.2 Results
Table˜3 summarizes our results. We find that SDPO outperforms GRPO across almost all runs, often leading to substantial improvements. SDPO learns notably faster than GRPO, performing close to 5 hours of GRPO training after only 1 hour of training with SDPO in several cases. SDPO achieves a particularly substantial improvement over GRPO on the Chemistry task, as is displayed in Figure 6 (left). With Olmo3-7B-Instruct, SDPO achieves the 5h GRPO accuracy in 30 minutes of wall-clock training time, a $10\times$ speedup. Moreover, SDPO’s 5h accuracy is more than 2020%-points higher than that of GRPO.
@@
Qwen3-8B 35.6 59.2 27.9 58.9 57.5
+ GRPO $54.7$ $60.0$ $63.8$ $72.7$ $34.3$ $51.8$ 74.3 $77.1$ $64.9$ $67.7$
+ GRPO (on-policy) 54.2 69.6 63.6 63.6 44.4 44.4 73.9 74.1 60.2 65.7
+ SDPO (on-policy) 60.0 70.1 66.6 75.6 51.5 52.9 72.1 78.4 68.0 68.5
Olmo3-7B-Instruct 18.8 37.7 18.1 36.7 39.3
+ GRPO 32.7 $46.8$ 55.3 $63.3$ 47.8 62.0 70.9 $75.0$ 56.4 65.0
+ GRPO (on-policy) 48.8 54.3 62.7 62.7 54.2 63.8 73.3 73.5 56.8 60.6
+ SDPO (on-policy) 59.2 76.8 59.9 66.1 56.1 $58.3$ 73.7 79.1 60.8 $62.1$
@@ 3.3 Self-distillation learns to reason concisely
We consistently observe that SDPO produces substantially shorter generations than GRPO while achieving higher accuracy. SDPO’s responses are more than $3\times$ shorter on average across tasks (cf. Table˜8 in Appendix˜D). On Chemistry with Olmo3-7B-Instruct, SDPO even achieves a $7\times$ reduction in response length relative to GRPO while maintaining higher accuracy (Figure 6 (right)). While recent progress in RLVR has demonstrated that scaling response length is a powerful driver of emergent reasoning capabilities (Jaech et al., 2024; Guo et al., 2025; Muennighoff et al., 2025), our results suggest that effective reasoning need not always be verbose. We find that SDPO improves the efficiency of reasoning.
Qualitatively, we observe that the longer responses from GRPO often stem from “superficial” reasoning rather than necessary cognitive steps. GRPO frequently generates filler phrases like “Hmm” and “Wait” or enters circular logical loops that repeat previous steps verbatim. Figure˜7 displays a representative example of this phenomenon. Remarkably, SDPO’s generations remain concise and avoid these superficial patterns. This may be explained by SDPO’s dense credit assignment, which assigns a specific advantage to each next-token prediction, leading to sparse advantages (cf. Figure˜21 in Appendix˜F). By improving the efficiency of reasoning, SDPO reduces inference generation time and demonstrates that reasoning performance can be improved by refining how the model reasons, not just how long it reasons.
@@ 4 Learning with Rich Environment Feedback
We use the Qwen3 (Yang et al., 2025) model family for our experiments, with Qwen3-8B as default unless otherwise specified. We report the average accuracy over 4 rollouts and use the same GRPO baseline as outlined in Section˜3.1.
@@ 4.1 Self-distillation benefits from stronger models
To answer this question, we perform a scaling study with different model sizes from the Qwen3 (Yang et al., 2025) family. As shown by extensive prior work, the ability to learn in-context increases with model size (e.g., Brown et al., 2020). As depicted in Figure˜8, SDPO significantly outperforms GRPO on larger models while only slightly improving over GRPO on smaller models. To determine whether SDPO can also underperform GRPO on a model weaker than Qwen3-0.6B, we performed an additional scaling study with Qwen2.5-Instruct (Qwen et al., 2024). While outperforming GRPO with Qwen2.5-7B and performing similarly with Qwen2.5-8B, we find that SDPO underperforms GRPO on Qwen2.5-1.5B, as seen in Figure˜17 in Appendix˜D.
@@ 4.2 Self-distillation performs dense credit assignment
Whereas GRPO assigns a constant advantage to each generated token, SDPO assigns an individual advantage to each possible next token along the generated sequence based on the agreement of student and teacher. At each position $t$ in the generated sequence $y$, there are $|\mathcal{V}|$ possible next tokens where $\mathcal{V}$ is the vocabulary. In distillation, this level is typically called the logit-level since it corresponds to the logits of the model. In practice, we approximate the full next-token distribution by the top-$K$ tokens, and as such, SDPO assigns $|y|\cdot K$ unique advantages per sequence. This is illustrated in Figure˜9 and allows SDPO to perform dense credit assignment.
@@ 4.4 On-policy self-distillation avoids catastrophic forgetting
Prior work has shown that a key benefit of on-policy algorithms, such as GRPO, is that models tend not to forget previously obtained capabilities (Shenfeld et al., 2026; Chen et al., 2025b; Lu & Thinking Machines Lab, 2025). This is practically desirable since it enables continual training pipelines where a model is trained sequentially on diverse tasks without the need to retrain from scratch. To evaluate forgetting, we test the final checkpoints of GRPO and SDPO on diverse holdout tasks: IFEval (Zhou et al., 2023), which tests the ability of a model to follow precise format instructions; ArenaHard-v2 (Li et al., 2025), which is an LLM-judged benchmark of real-world instruction-following prompts derived from LMArena (Chiang et al., 2024); and MMLU-Pro (Wang et al., 2024b), which tests broad multi-task knowledge and reasoning. As displayed in Table˜5, SDPO learns the new task while mitigating degradation of initial capabilities, overall achieving a better performance–forgetting tradeoff than GRPO.
@@
As an additional baseline, we consider training the student via supervised fine-tuning (SFT) on successful generations from the self-teacher (Scheurer et al., 2023; Dou et al., 2024). [^7] SFT on a teacher’s predictions is a standard off-policy distillation approach (Kim & Rush, 2016). This requires $2\times$ the generations of SDPO for the same number of steps, since we have to generate from both the student and the teacher. We report SFT on the successes of the self-teacher, which achieves a higher accuracy than also including initial successes from the student in the SFT data. As shown in Table˜5, SFT on the self-teacher significantly underperforms SDPO on LCBv6, while leading to worse forgetting of prior capabilities. This mirrors prior findings on the instability of off-policy imitation (see, e.g., Agarwal et al., 2024).

### v2

@@ 1 Introduction
Nevertheless, the dominant RL recipe for LLM post-training remains bottlenecked by credit assignment. Most current approaches operate in the setting of reinforcement learning with verifiable rewards (RLVR): given a question $x$, the model samples an answer $y\sim\pi_{\theta}(\cdot\mid x)$} and receives a scalar reward $r\in\mathbb{R}$, often binary (e.g., unit-tests pass/fail in code generation). Modern policy gradient RLVR methods such as Group Relative Policy Optimization (GRPO; Shao et al., 2024) estimate advantages from these sparse outcome rewards. Furthermore, when all rollouts in a group receive the same (often zero) reward, GRPO advantages collapse to zero and learning stalls. To overcome this sparsity, one might prefer distillation from a strong teacher (Guo et al., 2025; Yang et al., 2025a; Lu & Thinking Machines Lab, 2025; Guha et al., 2026), which provides dense, token-level supervision. However, strong teachers are often unavailable in online learning, where the goal is to raise the capability ceiling beyond existing models.
@@
  • Learning without rich feedback (§3): We evaluate standard RLVR environments that do not return any feedback beyond scalar rewards. Here, SDPO treats successful attempts sampled in the current batch as “feedback” for failed attempts on the same question. We perform training runs on scientific reasoning and tool use, starting with Qwen3-8B and Olmo3-7B-Instruct. We find that SDPO outperforms a strong GRPO baseline that integrates recent improvements: 70.2% vs. 66.6% final accuracy on aggregate. SDPO achieves higher accuracy with up to $11\times$ shorter generation lengths compared to GRPO, demonstrating that effective reasoning need not be verbose.
@@ 2 SDPO: Self-Distillation Policy Optimization
We can derive the SDPO gradient as follows (see Section˜A.1 for details):
@@
The gradient of $\mathcal{L}_{\mathrm{SDPO}}$ is
$\boldsymbol{\nabla}\mathcal{L}_{\mathrm{SDPO}}(\theta)=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\!\!\left[\sum_{t=1}^{|y|}\mathbb{E}_{\hat{y}_{t}\sim\pi_{\theta}(\cdot\mid x,y_{<t})}\!\!\left[\log\frac{\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})}{\pi_{\theta}(\hat{y}_{t}\mid x,f,y_{<t})}\cdot\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\right]\right].$ (2)
@@ 2.1 Comparison to RLVR
Note that the SDPO gradient is a (negated) logit-level policy gradient where the advantages are estimated using the self-teacher. [^2] See Section˜A.4 for a detailed comparison of the SDPO gradient to the standard policy gradient. We can therefore reuse standard RLVR implementations and simply swap out the advantages. Let $y_{i}$ be the $i$-th rollout from a rollout group of size $G$ for question $x$, then comparing GRPO and SDPO we have:
$A_{i,t}^{\mathrm{GRPO}}:=r_{i}-\mathrm{mean}\{r_{i}\}_{i=1}^{G}\;\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(constant in $t$)}},\quad A_{i,t}^{\mathrm{SDPO}}(\hat{y}_{i,t})=\log\frac{\pi_{\theta}(\hat{y}_{i,t}\mid x,f_{i},y_{i,<t})}{\pi_{\theta}(\hat{y}_{i,t}\mid x,y_{i,<t})}.$
The GRPO advantages are applied only to the sampled token $y_{i,t}$ and are constant within a rollout $y_{i}$. [^3] We use the GRPO (Shao et al., 2024) advantage without normalization (Liu et al., 2025b). In contrast, the SDPO advantages are zero only for tokens where student and teacher perfectly agree. The SDPO advantage is positive for tokens which are more likely under the teacher while being negative for tokens which are less likely under the teacher. Thus, SDPO can be seen as a direct extension of standard RLVR methods in two ways:
@@
This tight connection to RLVR methods also enables a straightforward extension of the SDPO gradient from Equation˜2 to off-policy data via PPO-style clipped importance sampling (Schulman et al., 2017), see Section˜A.4.
@@ 2.2 Compute time & memory
Naively computing the KL divergence between student and teacher requires holding full logits of both models in memory. To avoid this, we approximate the KL divergence in the SDPO loss by performing top-$K$ distillation (i.e., only computing the top-$K$ logits of the student and the corresponding logits of the teacher alongside a term capturing the tail probability; cf. Section˜A.3). With a reasonable choice of $K$ (e.g., K=100{K=100}), this avoids virtually any memory overhead while capturing most of the information.
@@ 2.3 Stability improvements
We find that two practical modifications significantly enhance the training stability of SDPO. First, we employ a regularized self-teacher, implemented either via an exponential moving average (EMA) of the student parameters or by interpolating the current teacher with the initial teacher (cf. Section˜A.2). As detailed later, both strategies effectively stabilize learning. Second, we adopt the symmetric Jensen-Shannon divergence for the distillation loss; this formulation has similarly been shown to improve stability in on-policy distillation from external teachers (Agarwal et al., 2024).
@@ 3 Learning without Rich Environment Feedback
Figure 6: Training progression of Olmo3-7B-Instruct on Chemistry. We report the average accuracy across 16 samples per question and a rolling average of response lengths over 5 steps. We report GRPO with the optimal hyperparameters for this model and task. We run each configuration for 3 seeds and report standard errors as shaded areas.
@@ 3.1 Experimental setting
  • Science Q&A (Chemistry, Physics, Biology, Materials science): Undergraduate-level scientific reasoning using reasoning subsets (L3) from SciKnowEval (Feng et al., 2024a).
@@
We perform a train-test split to test in-domain generalization. We use Qwen3-8B (Yang et al., 2025a) and Olmo3-7B-Instruct (Olmo et al., 2025) as initial checkpoints and report avg@16 relative to wall-clock training time, excluding initialization & validation.
@@
We compare SDPO to an improved variant of GRPO (Shao et al., 2024), which incorporates several recent modifications (Olmo et al., 2025; Khatri et al., 2026) such as asymmetric clipping (Yu et al., 2025), avoiding biased normalization (Liu et al., 2025b), and correcting for off-policy data when using efficient inference frameworks (Yao et al., 2025). We integrate these modifications into a GRPO implementation that represents a strong baseline, as detailed in Equation˜14 in Section˜A.4. GRPO enables off-policy training through PPO’s clipped importance weighting (Schulman et al., 2017). We additionally report the special case of on-policy GRPO (matching the hyperparameters of vanilla SDPO). For both baselines, we perform a hyperparameter sweep and report results for the models that achieve the highest validation performance across all target tasks. Hyperparameters and training details are provided in Appendix˜E. We use the verl library (Sheng et al., 2025) for fast multi-GPU training.
@@ 3.2 Results
Table˜3 summarizes our results. We find that SDPO outperforms GRPO across almost all runs, often leading to substantial improvements. SDPO learns notably faster than GRPO, performing close to 5 hours of GRPO training after only 1 hour of training with SDPO in several cases. SDPO achieves a particularly substantial improvement over GRPO on the Chemistry task, as is displayed in Figure 6 (left). With Olmo3-7B-Instruct, SDPO achieves the 5h GRPO accuracy in 50 minutes of wall-clock training time, a $6\times$ speedup. Moreover, SDPO’s 5h accuracy is more than 1010%-points higher than that of GRPO.
@@
Qwen3-8B 41.2 59.2 30.8 58.9 57.5
+ GRPO $65.9$ $74.5$ $63.8$ $72.7$ $35.1$ 59.9 74.3 $77.1$ $64.9$ $67.7$
+ GRPO (on-policy) 63.3 63.4 63.6 63.6 49.8 49.8 73.9 74.1 60.2 65.7
+ SDPO (on-policy) 73.2 80.9 66.6 75.6 50.6 56.8 72.1 78.4 68.0 68.5
Olmo3-7B-Instruct 22.8 37.7 16.2 36.7 39.3
+ GRPO 39.7 56.7 55.3 $63.3$ 35.6 55.8 70.9 $75.0$ 56.4 65.0
+ GRPO (on-policy) 51.4 57.5 62.7 62.7 49.8 49.8 73.3 73.5 56.8 60.6
+ SDPO (on-policy) 68.0 80.0 59.9 66.1 48.0 $52.8$ 73.7 79.1 60.8 $62.1$
@@ 3.3 Self-distillation learns to reason concisely
We consistently observe that SDPO produces substantially shorter generations than GRPO while achieving higher accuracy. SDPO’s responses are more than $3\times$ shorter on average across tasks (cf. Table˜8 in Appendix˜D). On Chemistry with Olmo3-7B-Instruct, SDPO even achieves an $11\times$ reduction in response length relative to GRPO while maintaining higher accuracy (Figure 6 (right)). While recent progress in RLVR has demonstrated that scaling response length is a powerful driver of emergent reasoning capabilities (Jaech et al., 2024; Guo et al., 2025; Muennighoff et al., 2025), our results suggest that effective reasoning need not always be verbose. We find that SDPO improves the efficiency of reasoning.
Qualitatively, we observe that the longer responses from GRPO often stem from “superficial” reasoning rather than necessary analytical steps. GRPO frequently generates filler phrases like “Hmm” and “Wait” or enters circular logical loops that repeat previous steps verbatim. Figure˜7 displays a representative example of this phenomenon. Remarkably, SDPO’s generations remain concise and avoid these superficial patterns. This may be explained by SDPO’s dense credit assignment, which assigns a specific advantage to each next-token prediction, leading to sparse advantages (cf. Figure˜21 in Appendix˜F). By improving the efficiency of reasoning, SDPO reduces inference generation time and demonstrates that reasoning performance can be improved by refining how the model reasons, not just how long it reasons.
@@ 4 Learning with Rich Environment Feedback
We use the Qwen3 (Yang et al., 2025a) model family for our experiments, with Qwen3-8B as default unless otherwise specified. We report the average accuracy over 4 rollouts and use the same GRPO baseline as outlined in Section˜3.1.
@@ 4.1 Self-distillation benefits from stronger models
To answer this question, we perform a scaling study with different model sizes from the Qwen3 (Yang et al., 2025a) family. As shown by extensive prior work, the ability to learn in-context increases with model size (e.g., Brown et al., 2020). As depicted in Figure˜8, SDPO significantly outperforms GRPO on larger models while only slightly improving over GRPO on smaller models. To determine whether SDPO can also underperform GRPO on a model weaker than Qwen3-0.6B, we performed an additional scaling study with Qwen2.5-Instruct (Qwen et al., 2024). While outperforming GRPO with Qwen2.5-7B and performing similarly with Qwen2.5-8B, we find that SDPO underperforms GRPO on Qwen2.5-1.5B, as seen in Figure˜17 in Appendix˜D.
@@ 4.2 Self-distillation performs dense credit assignment
Whereas GRPO assigns a constant advantage to each generated token, SDPO assigns an individual advantage to each possible next token along the generated sequence based on the agreement of student and teacher. At each position $t$ in the generated sequence $y$, there are $|\mathcal{V}|$ possible next tokens where $\mathcal{V}$ is the vocabulary. In distillation, this level is typically called the logit-level since it corresponds to the logits of the model. In practice, we approximate the full next-token distribution by the top-$K$ tokens plus the tail, and as such, SDPO assigns $|y|\cdot(K+1)$ unique advantages per sequence. This is illustrated in Figure˜9 and allows SDPO to perform dense credit assignment.
@@ 4.4 On-policy self-distillation avoids catastrophic forgetting
Prior work has shown that a key benefit of on-policy algorithms, such as GRPO, is that models tend not to forget previously obtained capabilities (Shenfeld et al., 2026b; Chen et al., 2025b; Lu & Thinking Machines Lab, 2025). This is practically desirable since it enables continual training pipelines where a model is trained sequentially on diverse tasks without the need to retrain from scratch. To evaluate forgetting, we test the final checkpoints of GRPO and SDPO on diverse holdout tasks: IFEval (Zhou et al., 2023), which tests the ability of a model to follow precise format instructions; ArenaHard-v2 (Li et al., 2025a), which is an LLM-judged benchmark of real-world instruction-following prompts derived from LMArena (Chiang et al., 2024); and MMLU-Pro (Wang et al., 2024b), which tests broad multi-task knowledge and reasoning. As displayed in Table˜5, SDPO learns the new task while mitigating degradation of initial capabilities, overall achieving a better performance–forgetting tradeoff than GRPO.
@@
As an additional baseline, we consider training the student via supervised fine-tuning (SFT) on successful generations from the self-teacher (Scheurer et al., 2023; Dou et al., 2024; Zhou et al., 2025). [^7] SFT on a teacher’s predictions is a standard off-policy distillation approach (Kim & Rush, 2016). This requires $2\times$ the generations of SDPO for the same number of steps, since we have to generate from both the student and the teacher. We report SFT on the successes of the self-teacher, which achieves a higher accuracy than also including initial successes from the student in the SFT data. As shown in Table˜5, SFT on the self-teacher significantly underperforms SDPO on LCBv6, while leading to worse forgetting of prior capabilities. This mirrors prior findings on the instability of off-policy imitation (see, e.g., Agarwal et al., 2024).